# Machine learning for advancing low-temperature plasma modeling and simulation

### v1

@@ 4 Concluding remarks
Open access to literature, databases, research data and codes, as well as data-driven models is foreseen to be an integral part in these efforts.[220, 221, 222, 63] Sharing information following the findable, accessible, interoperable, and reusable (FAIR) principles[223] would allow researchers independent of status and financial budget to develop the field. We believe this may enable vital new pathways for research on LTP processing.
@@
Despite being generally true for science and technology, it may be envisioned that ML and data-driven approaches to LTP modeling and simulation may not only support the analysis of known unknowns. In contrast, due to its inherent propensity to spotlight hidden patterns in data, it may encourage discovery of unknown unknowns. This aspect is argued to offer unprecedented potential for advances in plasma science and technology and beyond.

### v2

@@ 4 Concluding remarks
Open access to literature, databases, research data and codes, as well as data-driven models are foreseen to be an integral part in these efforts.[224, 225, 226, 63] Sharing information following the findable, accessible, interoperable, and reusable (FAIR) principles[227] would allow researchers independent of status and financial budget to develop the field. We believe this may enable vital new pathways for research on LTP processing.
@@
Despite being generally true for science and technology, it may be envisioned that scientific ML and data-driven approaches to LTP modeling and simulation may not only support the analysis of known unknowns. In contrast, due to its inherent propensity to spotlight hidden patterns in data, it may encourage discovery of unknown unknowns. This aspect is argued to offer unprecedented potential for advances in plasma science and technology and beyond.